# Shear viscosity of quark-gluon plasma at finite temperature and chemical potential and QCD phase transitions

### v1

@@ III.2 Shear viscosity
We plot in Fig. 9 the value of $\eta/s$ on the boundaries of the chiral phase transition of $u,d$ quarks. Fig. 9(a), (b) and (c) present the behavior of $\eta/s$ as functions of temperature, baryon chemical potential and baryon density, respectively. Since the relation between $\rho_{B}$ and $\mu_{B}$ on the crossover phase transition line near the critical region is nonmonotonic as indicated in Fig. 2 and Fig. 3, Fig. 9(c) shows a similar structure for $\eta/s$. For the first-order phase transition, Fig. 9(a) also indicates that $\eta/s$ in the chiral restored QGP phase is smaller than that in the chiral breaking phase in the region not far away from the critical region. This situation reverses as the temperature decreases (correspondingly, the chemical potential increases as shown in Fig. 9(b)). Such a feature reflects that the value of $\eta/s$ at intermediate chemical potential and temperature are affected by the competition between the temperature, chemical potential (density) and QCD phase transition.
@@
Figure 10: $\eta/s$ along the different isentropic trajectories. The isentropic trajectories are plotted in Fig. 1. The solid triangles (dots) mark the temperature of chiral phase transition in the crossover region ( of the higher-density side in the first-order region).
Finally, we demonstrate in Fig. 10 the behavior of $\eta/s$ along the different isentropic trajectories plotted in Fig. 1. It can be seen that $\eta/s$ has a minimum on each isentropic trajectory. The value of $\eta/s$ changes slowly for larger $s/\rho_{B}$ except for the region near the chiral phase transition. This situation corresponds to the fireball generated in high energy collision. For the case of smaller $s/\rho_{B}$, the rapid change of $\eta/s$ occurs near the first-order phase transition with the associated spinodal structure. The numerical results demonstrate that the minimum value of $\eta/s$ on each curve appear near the location of the first-order phase transition, as indicated with the solid dots in Fig. 10. The solid dots mark the boundary of the first-order phase transition at the high-density side. It is helpful to understood the result in conjunction with QCD phase diagrams shown in Fig. 2.
@@
In this work, we explored the shear viscosity of QGP in the full QCD phase diagram within the framework of kinetic theory with the relaxation time approximation. The flavor dependent relaxation time is derived according to the $2\leftrightarrow 2$ elastic scattering between quasiparticles including the $u,d,s$ quarks and their antiparticles. The temperature and chemical potential dependent masses of quasiparticles are calculated in the PNJL model. The masses of exchanged mesons are derived with the random phase approximation.
The numerical results show that, for small chemical potential, the relaxation time for each kind of quark and its antiparticle has a minimum near the Mott dissociations of mesons in the high-temperature side. The relaxation time increases rapidly below the chiral crossover phase transition line. Correspondingly, from high temperature to lower one the value of $\eta/s$ has a concave behavior and increase rapidly in the lower-temperature side of the crossover phase transition. The minimum value of $\eta/s$ at $\mu_{B}=0$ is about 0.1, close to the KSS limit and the value extracted from the experimental data.
At large chemical potential (high density), the behavior of $\eta/s$ in the QGP phase is mainly dominated by temperature. The value of $\eta/s$ of dense QGP is greatly enhanced with the decline of temperature. $\eta/s$ is about 1 when approaching to the critical region from the high-temperature side. The situation becomes relatively complicated for the first-order phase transition. The value of $\eta/s$ is influenced by the competition between temperature, density effect and QCD phase transition. The evolving behaviors of $\eta/s$ in the QCD phase diagram can be referenced by hydrodynamics to simulate the QCD phase transition.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grant No. 11875213 and Natural Science Basic Research Plan in Shaanxi Province of China (Program No. 2024JC-YBMS-018).

### v2

@@ III.3 Shear viscosity under different conditions
We plot in Fig. 12 the value of $\eta/s$ on the boundaries of the chiral phase transition of $u,d$ quarks. Fig. 12(a), (b) and (c) present the behavior of $\eta/s$ as functions of temperature, baryon chemical potential and net baryon density, respectively. Since the relation between $\rho_{B}$ and $\mu_{B}$ on the crossover phase transition line near the critical region is nonmonotonic as indicated in Fig. 2 and Fig. 3, Fig. 12(c) shows a similar structure of $\eta/s$. For the first-order phase transition, Fig. 12(a) also indicates that $\eta/s$ in the chiral restored QGP phase is smaller than that in the chiral breaking phase in the region not far away from the critical region. This situation reverses as the temperature decreases (correspondingly, the chemical potential increases as shown in Fig. 12(b)). Such a feature reflects that the value of $\eta/s$ in the two phase of the chiral first-order phase transition are affected by temperature, chemical potential (density) and QCD phase transition together.
@@
Figure 13: $\eta/s$ on different isentropic trajectories in the PNJL model. The isentropic trajectories are plotted in Fig. 1. The dot, solid triangles and hollow triangles represent the $\eta/s$ at the points where these isentropic lines intersect the high-density side of the first-order phase transition line, chiral crossover phase transition line and pion Mott transition line, respectively.
We demonstrate in Fig. 13 the behavior of $\eta/s$ along the isentropic trajectories as plotted in Fig. 1 and Fig. 2. The dot, solid triangles and hollow triangles represent the $\eta/s$ at the points where these isentropic line intersect the high-density side of the first-order phase transition line, chiral crossover phase transition line and pion Mott transition line, respectively. It is helpful to understand the result in conjunction with QCD phase diagrams shown in Fig. 1 and Fig. 2. It can be seen in Fig. 13 that $\eta/s$ has a minimum on each isentropic trajectory. For the case of larger $s/\rho_{B}$, the isentropic lines, as shown in Fig. 1, are almost parallel to the temperature axis. Thus, it is straightforward to see that the behavior of the corresponding shear viscosity with growing temperature is similar to that with a fixed chemical potential (in the region of small chemical potential). For the case of smaller $s/\rho_{B}$, the isentropic curves, as shown in Fig. 2, traverse the first-order phase transition from low to high density, including the low-density chiral symmetry broken phase and the high-density chiral symmetry restored phase (with metastable and unstable phases existing between them). Due to the spinodal structure inherent in the first-order phase transition, the isentropic curve exhibits a complex trajectory in the $T-\mu_{B}$ phase diagram, as illustrated in Fig. 1. For $s/\rho_{B}=3,4,5$, the value of $\eta/s$ changes rapidly near the first-order phase transition with the associated spinodal structure. The numerical results show that the minimum value of $\eta/s$ on each curve occurs near the Mott transition, which is close to the high-density boundary of the first-order phase transition, as indicated by the solid dots in Fig. 10.
III.4 Electric conductivity of quark matter
The electrical conductivity ($\sigma$) is a key parameter characterizing the charge transport properties of QGP. It directly influences the electromagnetic response and evolution behavior of the plasma, and is related to physical issues such as phase transition dynamics and electromagnetic field evolution. In this section, we will explore the electrical conductivity of quark matter at finite temperature and density, in particular its behavior near the Mott and chiral phase transition with the inherent spinodal structure.
Figure 14: Electric conductivity to temperature ratio $\sigma/T$ as function of temperature at different chemical potential $\mu_{B}=0$, $400$, $800$, $873$ and 12001200\, MeV in the PNJL model. The dot, solid triangles and hollow triangles represent the $\sigma/T$ at the points where these $\mu_{B}(T)$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line, respectively.
We first display in Fig. 14 the ratio of the electrical conductivity to temperature, $\sigma/T$, as a function of temperature for different baryon chemical potentials. At small chemical potentials, a distinct minimum in $\sigma/T$ appears near the Mott transition, beyond which $\sigma/T$ is observed to increase with temperature. This trend is consistent with the lattice QCD calculations [100, 101, 102, 103]. Along with the chiral crossover phase transition, the value of $\sigma/T$ rises with decreasing temperature. This behavior aligns with hadronic model calculationss [104]. The characteristics of the electrical conductivity resemble those of the shear viscosity near the phase transition region. Specifically, both $\eta/s$ and $\sigma/T$ exhibit a concave-down structure with a pronounced minimum near the Mott transition, indicating that intense collisions and resonances hinder the transport of momentum and electric charge. At large chemical potentials, a significant enhancement of $\sigma/T$ is observed, which is primarily attributed to the increase in quark number density.
Figure 15: $\sigma/T$ as functions of net baryon density at different temperatures in the PNJL model. The solid triangles, dots, and hollow triangles represent the $\sigma/T$ at the points where these $T(\rho_{B})$ lines intersect the chiral corssover transition line, the first-order chiral phase transition line and pion Mott transition line, respectively. For T=50,75,100T=50,75,100\,MeV, the two dots on each isotherm correspond to respectively the high-density chiral restoration phase and low-density chiral breaking phase on the binodal line of the first-order phase transition.
In Fig. 15, we present the behavior of $\sigma/T$ as a function of net baryon density at different temperatures. Above the critical temperature, the temperature dependence of $\sigma/T$ is similar to that of $\eta/s$. Below the critical temperature, $\sigma/T$ gradually increases from low to high density. The two solid dots on each isotherm mark the boundaries of the first-order phase transition: the dot at lower density corresponds to the chiral symmetry broken phase, while the dot at higher density corresponds to the chiral symmetry restored phase. Evidently, for the first-order phase transition, the value of $\sigma/T$ in the chiral symmetry restored phase is lower than that in the chiral symmetry broken phase.
Figure 16: $\sigma/T$ as functions of baryon chemical potential at different temperatures in the PNJL model. The solid triangles, dots, and hollow triangles represent the $\sigma/T$ at the points where these $T(\mu_{B})$ lines intersect the chiral corssover transition line, the first-order chiral phase transition line and pion Mott transition line respectively. For T=50,75,100T=50,75,100\,MeV, the two dots on each curve correspond to respectively the high-density chiral restoration phase and low-density chiral breaking phase on the binodal line of the first-order phase transition.
We correlate the behavior of $\sigma/T$ with net baryon number density by projecting it onto the $T$-$\mu_{B}$ phase diagram in Fig. 16. The dashed curves on the isotherms delineate the metastable and unstable regions associated with the first-order phase transition. Fig. 15 and Fig. 16 reveal a significant enhancement of the electrical conductivity in the low-temperature, high-density regime (large $\mu_{B}$) relative to the high-temperature, low-density region.
Fig. 17 displays the value of $\sigma/T$ on the crossover transition line and the first-order phase transition boundaries. Along the chiral crossover phase transition line, $\sigma/T$ decreases monotonically as temperature drops. At the first-order transition, $\sigma/T$ is consistently larger in the chiral restored (high-density) phase than in the broken (low-density) phase. This trend differs from that of $\eta/s$ near the critical region in the first-order transition.
Figure 17: The value of $\sigma/T$ on the phase boundaries including the chiral corssover and first-order phase transitions in the PNJL model.
It is also interesting to calculate the dimensionless ratio of shear viscosity to electric conductivity, an interesting quantity which possibly has certain universal features [105, 106, 107, 108]. We present in Fig. 18 the ratio of $\eta/s$ to $\sigma/T$ as functions of temperature at different chemical potentials. It can be observed that the $(\eta/s)/(\sigma/T)$ increase rapidly with decreasing temperature near the Mott and chiral crossover phase transitions. Its value on the chiral crossover phase transition line ranges from approximately 12 to 18, which is consistent with the calculation in Ref. [109]. Different from the non-monotonic behaviors of $\eta/s$ and $\sigma/T$ with temperature in Fig. 6 and Fig. 14, $(\eta/s)/(\sigma/T)$ for each chemical potential always decreases monotonically with growing temperature. The value of $(\eta/s)/(\sigma/T)$ approaches 6.4 in the high-temperature regime. A similar result was derived in Ref. [110]. This behavior can be approximately understood as follows: at extremely high temperatures, with the chiral symmetry restoration of $u,d,s$ quarks, the differences between quark flavors (apart from the bare quark masses) become minimal. The system gradually approaches a state resembling a non-interacting ideal gas. In such a system, scale invariance dominates in the high-temperature limit, leading to a stabilization of the dimensionless $(\eta/s)/(\sigma/T)$.
Figure 18: The ratio of $\eta/s$ to $\sigma/T$ as a function of temperature at different chemical potentials in the PNJL model. The dot, solid triangles and hollow triangles represent the value at the points where the $\mu_{B}(T)$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line, respectively.
@@ III.4 Electric conductivity of quark matter
In this work, we explored the shear viscosity and electric conductivity of quark matter at finite temperature and density within the framework of kinetic theory with the relaxation time approximation. The flavor dependent relaxation time is derived according to the $2\rightarrow 2$ elastic scattering between quasiparticles including the $u,d,s$ quarks and their antiparticles. The temperature and chemical potential dependent masses of quasiparticles are calculated in the 2+1 flavor PNJL model. The masses of exchanged mesons are derived with the random phase approximation.
We focus on investigating the characteristic behaviors of shear viscosity and electrical conductivity across the Mott transition, chiral crossover transformation, and first-order phase transition associated with a spinodal structure. It is found that the relaxation time is critical in the description of shear viscosity and electrical conductivity. The numerical results show that, at small chemical potential, the relaxation time for each species of quark (anti-quark) has a minimum near the meson Mott dissociation, and increases rapidly below the chiral crossover phase transition line. Correspondingly, both the $\eta/s$ and $\sigma/T$ exhibit a concave behavior near the pion Mott transition, and increase rapidly at the crossover phase transition with the decrease of temperature. The minimum value of $\eta/s$ at vanishing chemical potential is about 0.1, close to the KSS limit and the value extracted from the experimental data. At the CEP, the value of $\eta/s$ is above 1 within the mean-field approximation.
The calculation also indicates that, at large chemical potential (high baryon density), the behavior of $\eta/s$ in the QGP phase is mainly dominated by temperature. The value of $\eta/s$ of dense QGP is greatly enhanced with the decline of temperature. At intermediate temperature and chemical potential, the values of $\eta/s$ are influenced by the competition between the temperature, density effect and QCD phase transition.
Generally, the $\sigma/T$ exhibits characteristics similar to those of $\eta/s$ in the QCD phase diagram. Our calculations further show that the dimensionless ratio $(\eta/s)/(\sigma/T)$ decreases monotonically with temperature and approaches 6.4 in the high-temperature limit. This asymptotic value reveals that quark matter in the PNJL model resembles a non-interacting ideal gas in the high-temperature limit and exhibits approximate scale invariance.
In the present study, the calculations are performed within the mean-field approximation. A more accurate description of shear viscosity and electrical conductivity in the critical region would necessitate calculations beyond the mean field and incorporate critical dynamical effects.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grant No. 12475145 and Natural Science Basic Research Plan in Shaanxi Province of China (Program No. 2024JC-YBMS-018).